# CINNAMON: A hybrid approach to change point detection and parameter estimation in single-particle tracking data

### v1

@@ 1 Introduction
In this paper, we are going to present our contribution to the 2nd AnDi Challenge. CINNAMON (i.e., Change poInt detectioN aNd pArameter estiMation fOr aNomalous diffusion) is a hybrid method using a combination of neural networks, feature-based algorithms, and analytical methods. In the single-trajectory task within the competition, it first identifies the change points by making use of a neural network consisting of a DAIN layer for data standardization, several 1D convolutional layers, LSTM layers and an attention layer [51]. Then, in each segment, the anomalous exponent is estimated with a gradient boosting method based on the features designed for the first AnDi Challenge [52]. For the diffusion coefficient, an estimator from [53] is used. In the ensemble task, another neural network is applied to identify the model which is used to generate trajectories. Its architecture is very similar to that for change point detection, as it contains the same building blocks.
@@ 2.1 Datasets and ground truth
for any $t,s\geq 0$ and the generalized diffusion coefficient $K>0$. The parameter $H\in(0,1)$, often referred to as the Hurst exponent, directly relates to the anomalous diffusion exponent $\alpha$ through the equation $\alpha=2H$. When $H=\frac{1}{2}$, the process corresponds to standard Brownian motion, with independent increments and a linear mean-squared displacement. For $H<\frac{1}{2}$, the process exhibits subdiffusion, often related to crowded environments. It corresponds to a slower than linear MSD growth. In contrast, for $H>\frac{1}{2}$, FBM is super-diffusive, which implies that motion is persistent and MSD grows faster than a linear function.
@@ 3 Methods
Our contribution to the 2nd AnDi Challenge roots, at least to some extent, in our previous experiences. CINNAMON (Change poInt detectioN aNd pArameter estiMation fOr aNomalous diffusion) is a hybrid method using a combination of neural networks, feature-based algorithms, and analytical methods. Its details are presented in the remaining part of this section.
@@ 3.1.1 Change point detection
For a trajectory $X=(x_{1},\ldots,x_{t})$, the trained classifier $\varphi:\mathbb{R}^{d\times w}\rightarrow\left[0,1\right]$ analyzes sliding windows of length $w$, generating probabilities $p_{i}=\varphi\left(X\left[i,i+w\right)\right)$, which indicate the the presence of change points. The detection algorithm identifies segments where $p_{i}$ exceeds a threshold $\gamma$, with change points estimated at probability maxima within these segments. The method assumes that change points can be detected through local patterns in the trajectory data, with the window length $w$ determining the scale of these patterns.
@@ 3.1.3 Estimation of $K$
where $X(t)$ is a pure anomalous diffusion process with diffusion coefficient $K$ and exponent $\alpha$, $\xi(t)$ denotes noise, which is assumed to be independent of $X(t)$ and normally distributed with zero mean and variance $\sigma^{2}$.
@@
Similarly, the time-averaged MSD (TAMSD) is also notably changed. Due to the nature of this competition and our strategy, we have chosen the TAMSD approach to the estimation of the coefficient $K$. We define its estimator for a sample $(x_{1},x_{2},x_{3},\ldots,x_{N})$ measured in times $\Delta t,2\Delta t,\ldots,N\Delta t$ as
$\widehat{MSD}(n\Delta t)=\frac{1}{N-n}\sum_{k=1}^{N-n}\left(x_{k+n}-x_{k}\right)^{2}.$ (4)
In order to estimate coefficient $K$, we just consider taking the logarithm of both sides of Eq. (3), which yields:
@@
By employing the TAMSD approach, this method ensures a more robust estimation of $K$ in the presence of measurement noise compared to classical ensemble-averaged MSD methods. The choice of $\tau_{\min}$ and $\tau_{\max}$ is crucial, as discussed in [53], since it defines the region where noise effects are minimized while maintaining a sufficient number of data points for a reliable estimation. In our implementation, we set $\tau_{\min}=1$ and $\tau_{\max}=10$ to balance statistical accuracy and noise minimization. However, a more refined selection of both $\tau_{\min}$ and $\tau_{\max}$ could likely improve the robustness of the estimation.
@@ 3.1.4 Classification of segments
Classification of segments of the trajectories turned out to be the most challenging part of the single-trajectory task. We tested several approaches, including the one that we proposed for the 1st AnDi Challenge [52] and finally ended up with a very simple procedure which classifies the segments based on the estimated values of $\alpha$ and $K$. Moreover, we completely neglected the ’’confined‘‘ state in the procedure (see Ref. [46] for explanation). Taking it into account would make the method more complicated with practically no impact on its accuracy.
Our procedure is shown in Fig. 6. If for a given segment the estimated $\alpha$ was large, the movement of the molecule in that segment was classified as directed. If, on the other hand, both $\alpha$ and $K$ were small, we considered the molecule to be immobile. In the remaining cases, we considered the movement of the molecule to be free.
@@ 3.1.5 Additional features from TDA
The engineering of the features for the 1st AnDi Challenge (see Ref. [52] and A) was guided by physical intuition and the statistical properties of the diffusive models under consideration. Since the models in the second challenge differ from the previous ones, we decided to further enrich the feature set by adding characteristics derived from topological data analysis (TDA).
It is a powerful framework that extracts shape-based features from complex data structures, providing insights into the global and local topological properties of a dataset [64, 65]. One of the key tools in TDA is the persistent homology, which characterizes multi-scale topological features by tracking the birth and death of topological structures, such as connected components and loops, across different scales [66].
Euler characteristic (EC) is one of the simplest topological invariants. In a nutshell, it characterizes geometrical features of a topological space defined by a data object. The EC is a scalar integer quantity defined as the alternating sum of the rank of topological bases [67]. Thus, the Euler characteristic $\chi$ of a cell complex $X$ may be written as
$\chi(X)=\sum_{n}(-1)^{n}k_{n},$ (9)
where $k_{n}$ denotes the number of elements (cells) of dimension $n$ in $X$. In case of a 2D manifold for instance, $\chi$ is nothing but the difference between the number of connected components and the number of holes.
By continuously changing a parameter (e.g., a threshold, radius or scale) and tracking how the EC evolves, one obtains an Euler characteristic profile (ECP), which provides a summary of the shape of the data [68]. When applied to particle trajectories, ECP can reveal structural patterns that are not captured by traditional statistical measures, thereby offering a complementary perspective on the nature of the diffusive process.
We will use Euler characteristics computed for complexes generated with the Vietoris-Rips filtration combined with the order of points within the trajectory (points that occur later in trajectory, occur later in filtration) as additional features for our regressor [68]. Here, a filtration is understood as a function that describes a dataset at different scales/resolutions. The Vietoris-Rips one is explained in Fig. 7.
Figure 7: An illustration of the Vietoris-Rips filtration combined with the order of points within a trajectory. In a given step, we consider only points from the trajectory up to time T. At each of these points, we place a circle of radius $\epsilon$. If the circles corresponding to two points from the trajectory intersect, we connect these points by an edge (e.g. left plot in the middle row). If the circles corresponding to 3 points intersect at a single point, we add a triangle based on those points (upper right plot). In the plots, trajectory points are marked in blue and are connected by gray lines representing particle‘s transitions. These blue vertices, together with edges (marked in black) and triangles (marked in red), form a complex whose Euler characteristic we calculate.
@@ 3.2 Ensemble task
After the classification of single trajectories within the experiment, its final label was determined with a majority vote. To find a mapping between the states of the particles and their parameters $(\alpha,K)$, we took the estimates of the parameters of the single-trajectory task and group them together with the k-means algorithm [71]. If the model was identified as the multi-state model (see Sec. 2), the number of clusters was determined with the silhouette score. In case of a single-state model, only one cluster was taken into account. For the other models, the number of clusters was 2.
@@
Table 1: Comparison of the neural network architectures used in the single-trajectory (see Fig. 4) and the ensemble tasks. The biggest difference between the architectures are the Max_Pooling layers added after every convolution step in the ensemble task, as well as the additional Flatten layer. The shapes of the layers differ as well. Another difference, not shown in the table, is the last activation function – in the ensemble task a softmax function was used instead of a sigmoid. $B$ is the size of each data batch, $T$ is the length of the trajectory.
@@ 4.1 Change point detection
The training and validation loss values are presented in Figure 9. Four neural network classifier architectures were compared: the full model described in Sec. 3.1.1, as well as three simplified variants (CNN+LSTM, CNN-only, and LSTM-only). All architectures incorporate the DAIN layer for preprocessing.
Figure 9: Training and validation loss different variants of the neural network classifier. The full model, described in Sec. 3.1.1, is compared with three simplified versions (CNN+LSTM, CNN-only, and LSTM-only).
@@ 4.2 Parameters estimation
The results of the estimation of parameter $\alpha$ using the features of the first AnDi challenge are presented in Fig. 10.
Figure 10: Prediction of $\alpha$ with the gradient boosting regressor and the feature set from the 1st AnDi Challenge.
@@
For details of the feature, please refer to A.
Figure 11: Importance of the six most important features for estimating $\alpha$ using gradient boosting. Impact of the remaining features on estimation is minimal. The presented values reflect how often the features are used to split the data across all trees.
@@
Figure 12: Prediction of $K$ with the method from Sec. 3.1.3.
@@
Adding topological descriptors to the set of characteristics did not improve the results either. Their effect on the estimation was negligible. More details regarding the performance of different approaches can be seen in Tables 5-6.
R2 MSE
AnDi 1 features 0.006 0.923
TDA + AnDi 1 features 0.006 0.923
@@
MSE R2
Estimator 0.010 0.864
AnDi 1 features 0.008 0.973
TDA + AnDi 1 features 0.008 0.973
@@ Appendix A Feature set for gradient boosting regression
$A=-\log\left(1-\frac{(\lambda_{1}-\lambda_{2})^{2}}{2(\lambda_{1}+\lambda_{2})}\right),$ (12)
where $\lambda_{1}$ and $\lambda_{2}$ are the principle radii of gyration, i.e. the eigenvalues of the tensor $\mathbf{T}$ [73].
@@
    where $R_{t}$ is calculated as $R_{t}=\max_{1\leq s\leq t}[X_{s}-\frac{s}{t}X_{t}]-\min_{1\leq s\leq t}[X_{s}-\frac{s}{t}X_{t}]$ and $S_{t}$ is standard deviation of process $X_{t}$.
@@
$MW=\frac{1}{2(N+1)}\sum_{t=0}^{N}\left|\textrm{sgn}\left(\overline{X}_{i+1}^{(m)}-\overline{X}_{i}^{(m)}\right)-\textrm{sgn}\left(\overline{X}_{i+1}^{(m)}-\overline{X}_{i}^{(m)}\right)\right|,$ (32)

### v2

@@ 1 Introduction
In this paper, we are going to present our contribution to the 2nd AnDi Challenge. CINNAMON (i.e. Change poInt detectioN aNd pArameter estiMation fOr aNomalous diffusion) is a hybrid method using a combination of neural networks, feature-based algorithms, and analytical methods. In the single-trajectory task within the competition, it first identifies the change points by making use of a neural network consisting of a DAIN layer for data standardization, several 1D convolutional layers, LSTM layers and an attention layer [51]. Then, in each segment, the anomalous exponent is estimated with a gradient boosting method based on the features designed for the first AnDi Challenge [52]. For the diffusion coefficient, an estimator from [53] is used. In the ensemble task, another neural network is applied to identify the model which is used to generate trajectories. Its architecture is very similar to that for change point detection, as it contains the same building blocks.
@@ 2.1 Datasets and ground truth
for any time $t,s\geq 0$ and the generalized diffusion coefficient $K>0$ with units $[K]=\frac{\mathrm{length}^{2}}{\mathrm{time}^{2H}}$. The parameter $H\in(0,1)$, often referred to as the Hurst exponent, directly relates to the anomalous diffusion exponent $\alpha$ through the equation $\alpha=2H$. When $H=\frac{1}{2}$, the process corresponds to standard Brownian motion, with independent increments and a linear mean-squared displacement. For $H<\frac{1}{2}$, the process exhibits subdiffusion, often related to crowded environments. It corresponds to a slower than linear MSD growth. In contrast, for $H>\frac{1}{2}$, FBM is super-diffusive, which implies that motion is persistent and MSD grows faster than a linear function.
@@ 3 Methods
Our contribution to the 2nd AnDi Challenge roots, at least to some extent, in our previous experiences. CINNAMON (Change poInt detectioN aNd pArameter estiMation fOr aNomalous diffusion) is a hybrid method using a combination of neural networks, feature-based algorithms, and analytical methods. Its details are presented in the remaining part of this section. The implementation of the method may be found on Github [55] .
@@ 3.1.1 Change point detection
The weights and biases of the neural network were initialized with the popular Glorot method [62]. The model was trained over 10 epochs using the Adam optimizer [63] with a learning rate equal to 0.001 and a batch size of 32. Training and validation metrics demonstrated consistent performance improvement with no signs of overfitting.
@@
For a trajectory $X=(x_{1},\ldots,x_{t})$, the trained classifier $\varphi:\mathbb{R}^{d\times T}\rightarrow\left[0,1\right]$ analyzes sliding windows of length $T$, generating probabilities $p_{i}=\varphi\left(X\left[i,i+T\right)\right)$, which indicate the likelihood of a change point occurring (see Fig. 5). The detection algorithm identifies segments where $p_{i}$ exceeds a threshold $\gamma$ ($\gamma=0.95$ was used throughout this work). Within each segment, the precise location of the change point is determined by finding the position where the probability reaches its maximum value. This approach eliminates the issue of detecting consecutive change points, as only a single point from each continuous segment above the threshold is classified as a change point. The method assumes that change points can be detected through local patterns in the trajectory data, with the window length $T$ determining the scale of these patterns. Experiments revealed that while classifiers trained on longer subsequences ($T=20$) showed better classification accuracy, a window length of $T=15$ provides the optimal balance between detection accuracy and the ability to identify closely spaced change points.
Figure 5: Explanation of the change point detection method. The classifier shown in Fig. 4 analyses a window of length $T=15$ sliding over a trajectory in order to generate probabilities $p_{i}$, which indicate the likelihood of a change point occuring. The algorithm identifies then segments where $p_{i}$ exceeds a threshold $\gamma=0.95$. Within each segment, the precise location of the change point is determined by finding the position where the probability reaches its maximum value. In the example presented in the diagram, the left change point would be identified correctly, the right one would be found as well but at wrong position and the method would fail to identify the point in the middle.
@@ 3.1.3 Estimation of $K$
where $X(t)$ is a pure anomalous diffusion process with diffusion coefficient $K$ and exponent $\alpha$, $\xi(t)$ denotes noise, which is assumed to be independent of $X(t)$ and normally distributed with zero mean and variance $\sigma^{2}$ for each $t$.
@@
The last expression may be rewritten as
$\langle\tilde{X}^{2}(n\Delta t)\rangle=2K(n\Delta t)^{\alpha}+\sigma^{2}=2K(n\Delta t)^{\alpha}\left(1+\frac{\sigma^{2}}{2K(n\Delta t)^{\alpha})}\right)$ (4)
Taking the logarithms of both sides yields
@@
In the last equation, the ensemble average MSD, $\langle\tilde{X}^{2}(n\Delta t)\rangle$, was replaced by the time-averaged MSD (TAMSD). The estimator of the latter for a sample $(x_{1},x_{2},x_{3},\ldots,x_{N})$ measured in times $\Delta t,2\Delta t,\ldots,N\Delta t$ is defined as
$\widehat{MSD}(n\Delta t)=\frac{1}{N-n}\sum_{k=1}^{N-n}\left(x_{k+n}-x_{k}\right)^{2},$ (8)
for $n=0,1,2,\ldots,N-1$.
@@
By employing the TAMSD approach, this method ensures a more robust estimation of $K$ in the presence of measurement noise compared to classical ensemble-averaged MSD methods. The choice of $\tau_{\min}$ and $\tau_{\max}$ is crucial, as discussed in [53], since it defines the region where noise effects are minimized while maintaining a sufficient number of data points for a reliable estimation. In our implementation, we set $\tau_{\min}=1$ and $\tau_{\max}=10$ to balance statistical accuracy and noise minimization. However, a more refined selection of both $\tau_{\min}$ and $\tau_{\max}$ could likely improve the robustness of the estimation. Alternatively, one could utilize non-linear regression of formula (5), with logatithm term, or its Taylor expansion, as in [53]. It could lead to a better $\alpha$ and $K$ estimation results, at a cost of higher computational cost. However, the trajectories available in the challenge were usually quite short, and such an approach would often lead to higher level of statistical errors.
@@ 3.1.4 Classification of segments
Classification of segments of the trajectories turned out to be the most challenging part of the single-trajectory task. We tested several approaches, including the one that we proposed for the 1st AnDi Challenge [52] (referred to as the GB-Classifier I) and finally ended up with a very simple procedure which classifies the segments based on the estimated values of $\alpha$ and $K$ (referred to as the naive method). Moreover, we completely neglected the ’’confined‘‘ state in the procedure (see Ref. [46] for explanation). Taking it into account would make the method more complicated with practically no impact on its accuracy.
The naive method is shown in Fig. 7. If for a given segment the estimated $\alpha$ was large, the movement of the molecule in that segment was classified as directed. If, on the other hand, both $\alpha$ and $K$ were small, we considered the molecule to be immobile. In the remaining cases, we considered the movement of the molecule to be free.
@@
In Table 7 (Sec. 4.3), a brief comparison of the naive method and the GB-Classifier I is presented. A third method, being a combination of those two, is also considered there (GB-Classifier II). In the latter approach, a gradient boosting classifier working with the features from Appendix A was used to distinguish the ’’confinement‘‘ state from the other ones. As for the rest of the states, it was processed further with the naive method form Fig. 7.
@@ 3.1.5 Additional features from TDA
The engineering of the features for the 1st AnDi Challenge (see Ref. [52] and A) was guided by physical intuition and the statistical properties of the diffusive models under consideration. Since the models in the second challenge differ from the previous ones, we decided to further enrich the feature set by adding characteristics derived from topological data analysis (TDA) [67, 68].
Euler characteristics computed for complexes generated with the Vietoris-Rips filtration [69] combined with the order of points within the trajectory (points that occur later in trajectory, occur later in filtration) have been used as additional features for our regressor.
@@ 3.2 Ensemble task
After the classification of single trajectories within the experiment, its final label was determined with a majority vote, .i.e. the most frequent label across all trajectories became the overall class of the experiment. To map between the diffusive states and their parameters $(\alpha,K)$, estimates of those parameters from the single-trajectory task were grouped together with the k-means algorithm [72]. No data normalization was carried out before clusterization. For the initialization of the centroids, the greedy k-means++ algorithms [73] was used. It was carried out only once for each clusterization. The number of groups was specified by the identified type of the model in most of the cases. For single-state models, only one cluster was considered. For all other models except the multi-state one, the number of clusters was set to 2. Since the multi-state model does not imply any particular number of clusters, in this case we performed the grouping sequentially for the number of clusters ranging from 2 to $2+N$, where $N$ is the number of $(\alpha,K)$ pairs. For each clustering attempt, the average silhouette score was calculated and the grouping with its highest value was selected. No specific remediation strategy was implemented for cases with low clustering performance.
@@
Table 1: Comparison of the neural network architectures used in the single-trajectory (see Fig. 4) and the ensemble tasks. The biggest difference between the architectures are the Max_Pooling layers added after every convolution step in the ensemble task, as well as the additional Flatten layer. The shapes of the layers differ as well. Another difference, not shown in the table, is the last activation function – in the ensemble task a softmax function was used instead of a sigmoid. $B$ is the size of each data batch ($B=32$ was used throughout this work), $T$ is the length of the trajectory.
@@ 4.1 Change point detection
The training and validation loss and accuracy values are presented in Figure 10 and Figure 9, respectively. Four neural network classifier architectures were compared: the full model described in Sec. 3.1.1, as well as three simplified variants (CNN+LSTM, CNN-only, and LSTM-only). All architectures incorporate the DAIN layer for preprocessing. The early plateau in the CNN-only model demonstrates that adding LSTM components significantly improves model performance, as the CNN+LSTM and the full models continue to decrease their loss throughout the 10 epochs. This confirms our hypothesis that capturing temporal dependencies is crucial for this task. Regarding the attention mechanism, while its contribution appears modest in the loss curves, accuracy metrics show that the full model achieves slightly better performance than CNN+LSTM (0.8140 accuracy vs. 0.8087). Even such a small improvement can be considered as valuable in the context of change-point detection. Additionally, the attention module adds minimal computational overhead to the model (194 640 vs. 190 415 parameters).
Figure 9: Training and validation accuracy for different variants of the neural network classifier. The full model, described in Sec. 3.1.1, is compared with three simplified versions (CNN+LSTM, CNN-only, and LSTM-only).
Figure 10: Training and validation loss for different variants of the neural network classifier. The full model, described in Sec. 3.1.1, is compared with three simplified versions (CNN+LSTM, CNN-only, and LSTM-only).
@@ 4.2 Parameters estimation
The results of the estimation of parameter $\alpha$ using the features of the first AnDi challenge are presented in Fig. 11. They were obtained for trajectories of length 100.
Figure 11: Prediction of $\alpha$ with the gradient boosting regressor and the feature set from the 1st AnDi Challenge. The error metrics for these estimates are presented in Table 5.
@@
For details regarding the features, please refer to A.
Figure 12: Importance of the six most important features for estimating $\alpha$ using gradient boosting. alpha is the standard estimation of the anomalous exponent, based on fitting the empirical TAMSD to the model (10). p_var_2 and p_var_3 correspond to exponents $\gamma^{p}$ fitted to the empirical p-variation (Eq. 26) with $p=2$ and $p=3$, respectively. max_ts is the mean maximal excursion (Eq. (18)). And finally, vac_lag_2 and vac_lag_1 are the velocity autocorrelation functions at lags 2 and 1 (Eq. (14)). Impact of the remaining features on estimation is minimal. The importances reflect how often the features are used to split the data across all trees.
@@
Figure 13: Prediction of $K$ with the method from Sec. 3.1.3. The error metrics for these estimates are presented in Table 6.
@@
Adding topological descriptors to the set of characteristics did not improve the results either. Their effect on the estimation was negligible. There are a few potential reasons for that outcome. First of all, since all of the diffusion models were based on the fractional Brownian motion, they might not generate trajectories with significant, persistent topological differences that the Euler characteristic can effectively distinguish. Secondly, the filtration we have chosen causes the initial parts of the trajectory to have a greater impact on the topological descriptor values than the final parts. As there is no reason to favor the initial points of the trajectory, this phenomenon may have had a negative impact on the quality of the proposed descriptor. More details regarding the performance of different approaches can be seen in Tables 5-6.
R2 MSE MAE
AnDi 1 features 0.973 0.008 0.069
TDA + AnDi 1 features 0.973 0.008 0.069
@@
R2 MSE MAE
Estimator 0.864 0.010 0.062
AnDi 1 features 0.923 0.006 0.051
TDA + AnDi 1 features 0.923 0.006 0.051
@@ 4.3 Diffusive states
One can of course ask the question how the predictive power of our method will change after taking that state into consideration. A comparison of our naive method with two simple feature-based classifiers is presented in Table 7. We see that the naive method indeed outperfoms the classifiers, which may indicate that the feature set designed for the previous edition of the AnDi Challenge may require far-reaching adjustments for the new type of data.
Naive GB-Classifier I GB-Classifier II
AVG 0.838 0.812 0.808
Weighted AVG 0.906 0.835 0.837
Table 7: F-1 score values for the classification of diffusive states, averaged over all experiments. Weighted averages are given in the second row (weights are the number of trajectories in each experiment). The definition of the classification methods is given in Sec. 3.1.4. The ’’naive‘‘ one was used throught this paper.
@@ Appendix A Feature set for gradient boosting regression
$A=-\log\left(1-\frac{(\lambda_{1}-\lambda_{2})^{2}}{2(\lambda_{1}+\lambda_{2})^{2}}\right),$ (12)
where $\lambda_{1}$ and $\lambda_{2}$ are the principle radii of gyration, i.e., the eigenvalues of the tensor $\mathbf{T}$ [75]. Higher values of $A$ indicate a greater tendency for a preferred direction.
@@
The closer it is to 1, the straighter the trajectory is. In contrast, $S$ close to 0 indicates no preferred direction on the trajectory.
@@
    where $R_{t}$ is calculated as $R_{t}=\max_{1\leq s\leq t}[X_{s}-\frac{s}{t}X_{t}]-\min_{1\leq s\leq t}[X_{s}-\frac{s}{t}X_{t}]$ and $S_{t}$ is standard deviation of process $X_{t}$, and $E$ is the mathematical expectation.
@@
$MW=\frac{1}{2(N+1)}\sum_{i=0}^{N}\left|\textrm{sgn}\left(\overline{X}_{i+1}^{(m)}-\overline{X}_{i}^{(m)}\right)-\textrm{sgn}\left(\overline{X}_{i+1}^{(m)}-\overline{X}_{i}^{(m)}\right)\right|,$ (32)